OpenStudy (highschoolmom2010):

The measure of an exterior angle of a regular polygon is given. Find the measure of an interior angle. Then find the number of sides.

OpenStudy (highschoolmom2010):

33. 36 35. 30

ganeshie8 (ganeshie8):

interior angle = 180 - exterior angle

OpenStudy (highschoolmom2010):

so for 33 I just do 180-36?

ganeshie8 (ganeshie8):

yes that gives the interior angle, how do u find number of sides ?

OpenStudy (highschoolmom2010):

im not 100% on that

ganeshie8 (ganeshie8):

yea use this :- exterior angles add up to 360

ganeshie8 (ganeshie8):

so if u have 'n' sides in a regular polygon, then n x exterior angle = 360

ganeshie8 (ganeshie8):

n x 36 = 360 n = 360/36 = 10 so its a 10 sided polygon

OpenStudy (highschoolmom2010):

oh then i do the (n-2)180 to find interiors correct?

ganeshie8 (ganeshie8):

does question asks to find "sum of interior angles" also ?

OpenStudy (highschoolmom2010):

it wants an interior angle

ganeshie8 (ganeshie8):

interior angle you have found already. interior angle = 180 - exterior angle

ganeshie8 (ganeshie8):

interior angle = 180 - 36 = 144

ganeshie8 (ganeshie8):

so for Q 33), the answer is this :- its a 10 sided polygon with an interior angle of 144

OpenStudy (highschoolmom2010):

so i do the same thing for 35 too right?

ganeshie8 (ganeshie8):

yes

OpenStudy (highschoolmom2010):

alright i think i can do this now :)

ganeshie8 (ganeshie8):

sounds great :)

OpenStudy (highschoolmom2010):

180-30=150

OpenStudy (highschoolmom2010):

150 is the interior

OpenStudy (highschoolmom2010):

n=360/30 n=12

OpenStudy (highschoolmom2010):

@ganeshie8 is that right?

ganeshie8 (ganeshie8):

thats correct ! gw :)